# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ 7. Bounding Mean Relevant Work in the $M/G/2$
In this section, we derive explicit lower bounds on mean relevant work in the $M/G/2$ using the Sep-ISQ-2 and AR-ISQ-2 systems. We start with the 2-server system because it demonstrates the core idea of our proof before we move on to the more general results concerning the $M/G/k$. In Section 10.1, we numerically compare our novel bounds on mean relevant work to the existing bounds in the literature.
We will use the following test functions to bound mean relevant work in the Sep-ISQ-2 and AR-ISQ-2 systems. We explain the intuition behind these test functions in Section 8.
@@
For any job size distribution $S$ such that $E[S^{3}]$ is finite, and any arrival rate $\lambda$, the expected total work in the ISQ-$2$ system is given by
@@
To do so, we first consider the constant-drift test function $g_{2}$, defined in Definition 7.1. We want to calculate the drift $G\circ g_{2}(w,i)$ for all possible values of $w$ and $i$. Recall from Sections 6 and 7.1 that the possible values of the speed $i$ are speeds $0,1/2,$ and $1$. When $i=0$, the work must be 00 by definition, while if $i>0$, the work must be positive.
When $w>0$ and $i=1$, $G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1.$ Arrivals cause a drift of $\lambda E[S]$, while work completion causes a drift of $-1$. When $w>0$ and $i=1/2$, applying Lemma 5.1, we have $G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1.$ The choice of the function $g_{2}$ ensures this drift property (See Section 8). When $w=0$ and $i=0$,
$G\circ g_{2}(0,0)=\dfrac{1}{2}\left(1+2\lambda\mathbb{E}[S]-\widetilde{S}(2\lambda))\right)=\lambda\mathbb{E}[S]-1+\left(\frac{3}{2}-\dfrac{1}{2}\widetilde{S}(2\lambda))\right).$
@@
$G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1+\left(\frac{3}{2}-\frac{1}{2}\widetilde{S}(2\lambda))\right)\mathbbm{1}_{\{i=0\}}.$
@@
Now we switch our focus to characterizing the mean work of the ISQ-$k$ system. We consider the affine-drift test function, $h_{2}$, defined in Definition 7.2. When $w>0$ and $i=1$ or $1/2$,$G\circ h_{2}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(-1+\lambda\mathbb{E}[S]).$ When $w=0$ we have, $G\circ h_{2}(0,0)=\lambda\mathbb{E}[S^{2}]+\mathbb{E}[S]+\frac{\widetilde{S}(2\lambda))-1}{2\lambda}.$ Together, we can summarize the drift over all states as
$G\circ h_{2}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(-1+\lambda\mathbb{E}[S])+\left(\mathbb{E}[S]+\dfrac{\widetilde{S}(2\lambda))-1}{2\lambda}\right)\mathbbm{1}_{\{i=0\}}.$
@@
Using Proposition 7.3, we can characterize the mean work of the truncated-ISQ-2 system in the Sep-ISQ-2 system. This provides us an exact mean relevant work formula for the Sep-ISQ-2 system. We therefore lower bound the mean relevant work of an $M/G/k$ system under arbitrary scheduling policy by Theorem 7.6. This constitutes the first lower bound of Theorem 4.1, namely Equation 4.1. Our characterization of mean relevant work in the Sep-ISQ-2 system is as follows:
@@
For an arbitrary threshold $x$, arbitrary job size distribution $S$, and arbitrary arrival rate $\lambda$, the expected relevant work in the separate-recycling-ISQ-2 system is exactly given by
$\mathbb{E}[W_{x}^{\text{sep-ISQ}}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}+\lambda(1-F_{S}(x))x^{2}.$ (7.5)
@@
The relevant work in the separate-recycling-ISQ-2 system is the sum of the total work in the truncated-ISQ-2 system and the relevant work in the separate $M/G/\infty$ system. Note that all work in the truncated-ISQ-2 system is relevant.
First, let us handle the separate $M/G/\infty$ system. The servers at the $M/G/\infty$ system will operate at a speed of $1/2$, and the arrival rate into this system is determined by $\lambda(1-F_{S}(x))$. Thus the expected relevant work at this separate server is $\lambda(1-F_{S}(x))x^{2}$.
[…]
$\mathbb{E}[W_{x}^{truncated{\text{-}}ISQ}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}.$
@@
To derive the ISQ-2 arbitrary-recycling lower bound, we will use the following modified ISQ-2 affine-drift test function $h_{2,x}$. Note that $h_{2,x}$ is a different function for each value of $x$, we provide intuition on how we derive $h_{2,x}$ in Section 8. Again, it is easy to verify that the modified ISQ-2 affine-drift test function satisfies the assumptions of Lemma 5.5.
@@
where $C_{2}(x,\lambda_{x}):=\frac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}.$ Here, the first argument of $C_{2}$ is the relevancy cutoff level $x$ of the arbitrary-recycling ISQ-2 system.
We are ready to state the main result for our AR-ISQ-2 lower bound, which provides us with another lower bound in mean relevant work in the $M/G/k$ by Theorem 7.6. This constitutes the second lower bound of Theorem 4.1, namely Equation 4.2.
@@
For an arbitrary threshold $x$, job size distribution $S$, arrival rate $\lambda$, and arbitrary recycling stream, expected relevant work in the arbitrary-recycling-ISQ-2 system is lower bounded by
$\mathbb{E}[W_{x}^{\text{arb-ISQ}}]\geq\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}\cdot\dfrac{1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}}{1-\rho_{x}}+\dfrac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}.$ (7.7)
@@
Let $\mathbb{E}_{r}[\cdot]$ denote the Palm expectation taken over the moments when the arbitrary arrival of jobs of size $x$ occur. By Lemma 5.1, the expected drift $G\circ h_{2,x}$ has two kinds of terms, the stochastic drift due to Poisson arrivals of rate $\lambda_{x}=\lambda F_{S}(x)$ and the recycling jumps due to the arbitrary arrivals of jobs of size $x$ with rate $\lambda(1-F_{S}(x))$.
Let $J_{x}(w,i):=h_{2,x}(w+x,\min(i+1/2,1))-h_{2,x}(w,i)$ denote the increase in the test function due to the arrival of a size-$x$ job. Then $\mathbb{E}_{r}[J(W_{x},I)]$ denotes the mean size of the recycling jump in stationarity. The expectation $\mathbb{E}_{r}[\cdot]$ can be interpreted as the Palm expectation associated with the random measure that records the cumulative number of recyclings. We define $\mathbb{E}_{r\mid w,i,a}[\cdot]$ to be the conditional Palm expectation in the immediate future of a specific state of the AR-ISQ-$k$ system. See, for example, [27], [32] and [2]. We start by applying Lemma 5.6 to find the drift in a specific state:
$G\circ h_{2,x}(w,i)=\lambda_{x}\mathbb{E}[S_{x}^{2}]+2w(-1+\lambda\mathbb{E}[S_{x}])+2C_{2}(x,\lambda_{x})\left(\mathbbm{1}_{\{i=0\}}+\frac{1}{2}\mathbbm{1}_{\{i=1/2\}}\right)+(\lambda-\lambda_{x})\mathbb{E}_{r\mid w,i,a}[J_{x}(w,i)]$
@@
$\quad\quad+(\lambda-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)].$
@@
$\mathbb{E}[W_{x}]=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda_{x})}{1-\rho_{x}}\left(\mathbb{P}(I^{r}=0)+\frac{1}{2}\mathbb{P}(I^{r}=1/2)\right)+\frac{(\lambda-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}.$ (7.8)
In the above equation, $\mathbb{P}(I^{r}=0)$ and $\mathbb{P}(I^{r}=1/2)$ are difficult to evaluate. However, we show that we can relate the two probabilities to the capped load $\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}$. To do so, we apply Lemma 5.6 to the test function $g(w,i)=w$ for all $(w,i)$, we have
@@
$\mathbb{E}[W_{x}]=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda_{x})(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})\mathbb{E}_{r}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}.$
Next, we want to lower bound $J_{x}(w,i)$ over all possible states in which a recycling could occur, and specifically the three cases $i=1,1/2,0$, respectively. In particular, we show that $J_{x}(w,i)$ is uniformly lower bounded by $x^{2}$ for any arbitrary job size distribution $S$ and arrival rate $\lambda$.
@@
    $J_{x}(0,0)=x^{2}+\frac{x}{\lambda_{x}}-\frac{1-e^{-2x\lambda_{x}}}{2\lambda_{x}^{2}}-\frac{C_{2}(x,\lambda_{x})(1-e^{-2x\lambda_{x}})}{\lambda_{x}}\geq x^{2}+\frac{x}{\lambda_{x}}-\frac{1-e^{-2x\lambda_{x}}}{2\lambda_{x}^{2}}-\frac{x(1-e^{-2x\lambda_{x}})}{2\lambda_{x}}.$
@@
    $2x\lambda_{x}-(1-e^{-2x\lambda_{x}})(1+x\lambda_{x})=(x\lambda_{x}-1)+(1+x\lambda_{x})e^{-2x\lambda_{x}}\geq 0$
@@
Therefore, because $J_{x}(w,i)$ is lower bounded by $x^{2}$ for all $w,i$, we know that $\mathbb{E}_{r}[J_{x}(W_{x},I)]$ is similarly lower bounded by $x^{2}$. We therefore have our desired lower bound on the mean relevant work in the arbitrary recycling ISQ-2 system:
$\mathbb{E}[W_{x}^{arb{\text{-}}ISQ}]\geq\frac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda)(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}.$

### v3

@@ 8. Bounding Mean Relevant Work in the $M/G/2$
In this section, we derive explicit lower bounds on mean relevant work in the $M/G/2$ using the Sep-ISQ-2 and Rec-ISQ-2 systems. We start with the 2-server system because it demonstrates the core idea of our proof before we move on to the more general results concerning the $M/G/k$. In Section 12, we numerically compare our novel bounds on mean relevant work to the existing bounds in the literature.
We will use the following test functions to bound mean relevant work in the Sep-ISQ-2 and Rec-ISQ-2 systems. We explain the intuition behind these test functions in Section 9.
@@
For any job size $S$ such that $E[S^{3}]$ is finite, and any arrival rate $\lambda$, the expected total work in the ISQ-$2$ system is given by
@@
To do so, we first consider the constant-drift test function $g_{2}$, defined in Definition 8.1. We want to calculate the drift $G\circ g_{2}(w,i)$ for all possible values of $w$ and $i$. Recall from Sections 6 and 8.1 that the possible values of the speed $i$ are speeds $0,1/2,$ and $1$. When $i=0$, the work must be 0 by definition, while if $i>0$, the work must be positive.
When $w>0$ and $i=1$, $G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1.$ Arrivals cause a drift of $\lambda E[S]$, while work completion causes a drift of 1-1. When $w>0$ and $i=1/2$, applying Lemma 7.1, we have $G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1.$ The choice of the function $g_{2}$ ensures this drift property (See Section 9). When $w=0$ and $i=0$,
$G\circ g_{2}(0,0)=\dfrac{1}{2}\left(1+2\lambda\mathbb{E}[S]-\widetilde{S}(2\lambda)\right)=\lambda\mathbb{E}[S]-1+\left(\frac{3}{2}-\dfrac{1}{2}\widetilde{S}(2\lambda)\right).$
@@
$G\circ g_{2}(w,i)=\lambda\mathbb{E}[S]-1+\left(\frac{3}{2}-\frac{1}{2}\widetilde{S}(2\lambda)\right)\mathbbm{1}_{\{i=0\}}.$
@@
[…]
$G\circ h_{2}(w,i)=\lambda\mathbb{E}[S^{2}]+2w(-1+\lambda\mathbb{E}[S])+\left(\mathbb{E}[S]+\dfrac{\widetilde{S}(2\lambda)-1}{2\lambda}\right)\mathbbm{1}_{\{i=0\}}.$
@@
Using Proposition 8.3, we can characterize the mean work of the truncated-ISQ-2 system in the Sep-ISQ-2 system. This provides us an exact mean relevant work formula for the Sep-ISQ-2 system. We therefore lower bound the mean relevant work of an $M/G/k$ system under arbitrary scheduling policy by Theorem 8.6. This constitutes the first novel relevant work lower bound of Theorem 4.1, namely Equation 4.4. Our characterization of mean relevant work in the Sep-ISQ-2 system is as follows:
@@
For an arbitrary threshold $x$, arbitrary job size $S$, and arbitrary arrival rate $\lambda$, the expected relevant work in the separate-ISQ-2 system is exactly given by
$\mathbb{E}[W_{x}^{\text{Sep-ISQ-}2}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}+\lambda(1-F_{S}(x))x^{2}.$ (8.5)
@@
The relevant work in the separate-ISQ-2 system is the sum of the total work in the truncated-ISQ-2 system and the relevant work in the separate $M/G/\infty$ system. Note that all work in the truncated-ISQ-2 system is relevant.
First, let us handle the separate $M/G/\infty$ system. The servers at the $M/G/\infty$ system operate at a speed of $1/2$, and the arrival rate into this system is determined by $\lambda(1-F_{S}(x))$. Thus the expected relevant work at this separate server is $\lambda(1-F_{S}(x))x^{2}$.
Jobs arrive into the truncated-ISQ-2 with a conditional size $S_{x}$, having a density given by $f_{S}(x)/F_{S}(x)$ and bounded third moment $\mathbb{E}[S_{x}^{3}]\leq x^{3}$. The arrival rate is $\lambda_{x}:=\lambda F_{S}(x)$. In particular, $\mathbb{E}[S_{x}]=\int_{0}^{x}sf(s)/F_{S}(x)ds$, $\mathbb{E}[S^{2}_{x}]=\int_{0}^{x}s^{2}f(s)/F_{S}(x)ds$, and $\widetilde{S}_{x}(2\lambda_{x})=\int_{0}^{x}e^{-2\lambda_{x}s}f_{S}(s)/F_{S}(x)ds$. By Proposition 8.3, we have the following formula for expected relevant work formula for the truncated-ISQ-2 system:
$\mathbb{E}[W_{x}^{{\text{truncated -ISQ-}}2}]=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\lambda_{x}\mathbb{E}[S_{x}])}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}.$
@@
To derive the Rec-ISQ-2 lower bound, we will use the following modified ISQ-2 affine-drift test function $h_{2,x}$. Note that $h_{2,x}$ is a different function for each value of $x$. We provide intuition on how we derive $h_{2,x}$ in Section 9. Again, it is easy to verify that the modified ISQ-2 affine-drift test function satisfies the assumptions of Lemma 7.5.
@@
where $C_{2}(x,\lambda_{x}):=\frac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}.$ Here, the first argument of $C_{2}$ is the relevancy cutoff level $x$ of the recycling ISQ-2 system.
We are ready to state the main result for our Rec-ISQ-2 lower bound, which provides us with another lower bound in mean relevant work in the $M/G/k$ by Theorem 8.6. This constitutes the second lower bound of Theorem 4.1, namely Equation 4.5.
@@
For an arbitrary threshold $x$, job size $S$, arrival rate $\lambda$, and arbitrary recycling stream, expected relevant work $\mathbb{E}[W_{x}^{\text{Rec-ISQ-2}}]$ in the Rec-ISQ-2 system is lower bounded by
$\mathbb{E}[W_{x}^{\text{Rec-ISQ-2-}L}]:=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}\cdot\dfrac{1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}}{1-\rho_{x}}+\dfrac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}.$ (8.7)
@@
Let $\mathbb{E}_{\mathfrak{r}}[\cdot]$ denote the Palm expectation taken over the moments when the arbitrary arrival of jobs of size $x$ occurs according to the recycling stream. By Lemma 7.1, the expected drift $G\circ h_{2,x}$ has two kinds of terms, the drift due to Poisson arrivals of rate $\lambda_{x}=\lambda F_{S}(x)$ and the recycling jumps due to the arbitrary arrivals of jobs of size $x$ with rate $\lambda(1-F_{S}(x))$.
Let $J_{x}(w,i):=h_{2,x}(w+x,\min(i+1/2,1))-h_{2,x}(w,i)$ denote the increase in the test function due to the arrival of a size-$x$ job. Then $\mathbb{E}_{\mathfrak{r}}[J(W_{x},I)]$ denotes the mean size of the recycling jump in stationarity. The expectation $\mathbb{E}_{\mathfrak{r}}[\cdot]$ can be interpreted as the Palm expectation associated with the random measure that records the cumulative number of recyclings. We define $\mathbb{E}_{\mathfrak{r}\mid w,i,r}[\cdot]$ to be the conditional Palm expectation in the immediate future of a specific state of the Rec-ISQ-$k$ system (Braverman2024, Miyazawa1994, and Scully2020). We start by applying Lemma 7.6 to find the drift in a specific state:
$G\circ h_{2,x}(w,i)=\lambda_{x}\mathbb{E}[S_{x}^{2}]+2w(-1+\lambda\mathbb{E}[S_{x}])+2C_{2}(x,\lambda_{x})\left(\mathbbm{1}_{\{i=0\}}+\frac{1}{2}\mathbbm{1}_{\{i=1/2\}}\right)+(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}\mid w,i,r}[J_{x}(w,i)]$
@@
$\quad\quad+(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)].$
@@
$\mathbb{E}[W_{x}]=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda_{x})}{1-\rho_{x}}\left(\mathbb{P}(I^{r}=0)+\frac{1}{2}\mathbb{P}(I^{r}=1/2)\right)+\frac{(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}.$ (8.8)
In the above equation, we write $I^{r}$ instead of $I$ to make explicit the dependence of $I$ on recycling arrivals, and to avoid confusion with the probability $\mathbb{P}(I=0)$ defined in Equation 8.4.
Both $\mathbb{P}(I^{r}=0)$ and $\mathbb{P}(I^{r}=1/2)$ are difficult to evaluate. However, we show that we can relate the two probabilities to the capped load $\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}$. To do so, we apply Lemma 7.6 to the test function $g(w,i)=w$ for all $(w,i)$, we have
@@
$\mathbb{E}[W_{x}]=\frac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda_{x})(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]}{2(1-\rho_{x})}.$
Next, we want to lower bound $J_{x}(w,i)$ over all possible states in which a recycling could occur, and specifically the three cases $i=1,1/2,0$, respectively. In particular, we show that $J_{x}(w,i)$ is uniformly lower bounded by $x^{2}$ for any arbitrary job size $S$ and arrival rate $\lambda$.
@@
    $J_{x}(0,0)$ $=x^{2}+\frac{x}{\lambda_{x}}-\frac{1-e^{-2x\lambda_{x}}}{2\lambda_{x}^{2}}-\frac{C_{2}(x,\lambda_{x})(1-e^{-2x\lambda_{x}})}{\lambda_{x}}$
    $\geq x^{2}+\frac{x}{\lambda_{x}}-\frac{1-e^{-2x\lambda_{x}}}{2\lambda_{x}^{2}}-\frac{x(1-e^{-2x\lambda_{x}})}{2\lambda_{x}}.$
@@
    $2x\lambda_{x}-(1-e^{-2x\lambda_{x}})(1+x\lambda_{x})=(x\lambda_{x}-1)+(1+x\lambda_{x})e^{-2x\lambda_{x}}\geq 0.$
@@
Therefore, because $J_{x}(w,i)$ is lower bounded by $x^{2}$ for all $w,i$, we know that $\mathbb{E}_{\mathfrak{r}}[J_{x}(W_{x},I)]$ is also lower bounded by $x^{2}$. We therefore have our desired lower bound on the mean relevant work in the Rec-ISQ-2 system:
$\mathbb{E}[W_{x}^{\text{Rec-ISQ-2}}]\geq\frac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{C_{2}(x,\lambda)(1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu})}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}:=\mathbb{E}[W_{x}^{\text{Rec-ISQ-2-}L}].$
8.1. Proof of Theorem 4.1
In this subsection, we prove Theorem 4.1.
Proof.
To prove Theorem 4.1 we combine results established in Section 5 and Section 8. We obtain Equation 4.1 by applying the WINE formula to Equations 5.2 and 5.3.
Equation 4.2 further incorporates $\mathbb{E}[W_{x}^{\text{Sep-ISQ-2}}]$. The fact that Sep-ISQ-2 lower bounds the relevant work of any $M/G/2$ system under arbitrary scheduling policy is proved in Theorem 6.2 and we derive its analytical expression given by Equation 4.4 in Theorem 8.4.
Finally, Equation 4.3 incorporates $\mathbb{E}[W_{x}^{\text{Rec-ISQ-2-}L}]$. The fact that there exists a Rec-ISQ-2 system lower bounding the relevant work of any $M/G/2$ system under arbitrary scheduling policy is proved in Theorem 6.3 and we derive a lower bound on $\mathbb{E}[W_{x}^{\text{Rec-ISQ-2}}]$ given by Equation 4.5 in Theorem 8.4.